# Evolutive sandpiles

### v2

@@ 5 Conclusions
For example, within the algebraic perspective of the sandpiles, known as sandpile group. The sum of two configurations ${c}$ and ${d}$ is performed entry by entry. The sandpile sum ${c}\oplus{d}$ of two configurations is defined as $s({c}+{d})$. A configuration ${c}$ is recurrent if there exists a non-zero configuration ${d}$ such that ${c}={c}\oplus{d}$. Recurrent configurations play a central role in the dynamics of the Abelian sandpile model since recurrent configurations together with the sandpile sum form an Abelian group known as sandpile group [15, Chapter 4]. The sandpile group of $G$ is denoted by $K(G)$. One of the interesting features of the sandpile group of connected graphs is that the order $|K(G)|$ is equal to the number $\tau(G)$ of spanning trees of the graph $G$. For the reader interested in the mathematics of the sandpiles, we recommend the book [15] or the survey [1]. Is it possible to obtain such algebraic structure on non-trivial evolutive graphs.
@@ Acknowledgement
The authors would like to thank Lothar Dirks for his helpful comments. The research of C.A. Alfaro is partially supported by the Sistema Nacional de Investigadores grant number 220797.

### v3

@@ 5 Conclusions
For example, within the algebraic perspective of the sandpiles, known as sandpile group. The sum of two configurations ${c}$ and ${d}$ is performed entry by entry. The sandpile sum ${c}\oplus{d}$ of two configurations is defined as $s({c}+{d})$. A configuration ${c}$ is recurrent if there exists a non-zero configuration ${d}$ such that ${c}={c}\oplus{d}$. Recurrent configurations play a central role in the dynamics of the Abelian sandpile model since recurrent configurations together with the sandpile sum form an Abelian group known as sandpile group [16, Chapter 4]. The sandpile group of $G$ is denoted by $K(G)$. One of the interesting features of the sandpile group of connected graphs is that the order $|K(G)|$ is equal to the number $\tau(G)$ of spanning trees of the graph $G$. For the reader interested in the mathematics of the sandpiles, we recommend the book [16] or the survey [1]. Is it possible to obtain such an algebraic structure on non-trivial evolutive graphs?
@@ Acknowledgement
The authors would like to thank Lothar Dirks for his helpful comments. The research of C.A. Alfaro is partially supported by the Sistema Nacional de Investigadores grant number 220797. The research of Ralihe R. Villagran was supported, in part, by the National Science Foundation through the DMS Award ##\##1808376 which is gratefully acknowledged.